1. The energy 𝑖, of an inductor with inductance 𝐿 is given by

𝑖= 1/2𝐿∫10 𝑑^2𝑒^βˆ’π‘‘ 𝑑𝑑

For 𝐿=(1 Γ—10βˆ’3)𝐻, Find 𝑖.


1
Expert's answer
2021-05-31T19:33:34-0400

Let us find the energy:


"𝑖= \frac{1}{2}𝐿\int_0^1 𝑡^2𝑒^−𝑡 𝑑𝑡=|u=t^2,\ dv=e^{-t}dt,\ du=2tdt,\ v=-e^{-t}|=
\frac{1}{2}𝐿(-t^2e^{-t}|_0^1+2\int_0^1te^{-t}dt)=\frac{1}{2}𝐿(-e^{-1}+2\int_0^1te^{-t}dt)=|u=t, dv=e^{-t}dt, du=dt, v=-e^{-t}|=\frac{1}{2}𝐿(-e^{-1}+2(-te^{-t}|_0^1+\int_0^1e^{-t}dt))=
-e^{-t}|=\frac{1}{2}𝐿(-e^{-1}+2(-e^{-1}-e^{-t}|_0^1))=
\frac{1}{2}𝐿(-e^{-1}+2(-e^{-1}-e^{-1}+1))=
\frac{1}{2}𝐿(2-5e^{-1})"


For "\ud835\udc3f=(1 \u00d710^{\u22123})H" we have "i=\frac{1}{2}(2-5e^{-1})10^{-3}H."